OpenStudy (anonymous):

can someone simplfy this one 1/3x+9 - 2/2x3=2

OpenStudy (dumbcow):

haha this time there is an equals sign so you can solve for x....but simplifying it will help is this it? \[\frac{1}{3x+9} +\frac{2}{2x^{3}} = 2\]

OpenStudy (dumbcow):

oh i meant "-"

OpenStudy (dumbcow):

but denominators are correct

OpenStudy (dumbcow):

yes that is wrong

OpenStudy (anonymous):

ok, so r u saying the answer is neg 2

OpenStudy (dumbcow):

no, i haven't given an answer yet...i want to make sure i have the right equation first :)

OpenStudy (dumbcow):

i don't think that is right...have you learned how to solve for roots of cubic function yet?

OpenStudy (anonymous):

yep, I have been solving problems like this all day, I think I just had a brain fart. LOL

OpenStudy (dumbcow):

ok well combine the fractions on left side Note: the 2's cancel in second fraction \[\frac{x^{3} -3x-9}{x^{3}(3x+9)} = 2\]

OpenStudy (dumbcow):

multiply denominator over \[x^{3} -3x-9 = 2x^{3}(3x+9)\] move everything to 1 side \[6x^{4}+17x^{3}+3x+9 = 0\] then i would graph the function to get one real root, then factor it out to get the rest... this really does not seem right for what you have been doing ??

OpenStudy (anonymous):

no problem, thanks anyway. let me try someone else. I have to get this done

OpenStudy (dumbcow):

no what i mean is i think your question isn't clear... here is solution , does this match any possible answers? http://www.wolframalpha.com/input/?i=6x^4%2B17x^3%2B3x%2B9%3D0